OpenStudy (anonymous):

Find the dot product, a • b. a = 6i + 5j, b = -5i + 4j

OpenStudy (michele_laino):

are a, b vectors?

OpenStudy (anonymous):

The answers are: 50 -10 <1, 9> <-30, 20> Would it be -10 or <-30, 20>? I forget if you leave it or add it together.

OpenStudy (michele_laino):

you have to multiply component times component, namely: \[a*b=(6*-5)+(5*4)=-30+20=-10\]

OpenStudy (anonymous):

Thanks! @Michele_Laino

OpenStudy (michele_laino):

since i*j=0, i*i=j*j=1

OpenStudy (michele_laino):

Thanks!